OpenStudy (anonymous):

The sum of two consecutive even integers is 814.

OpenStudy (imstuck):

What is the question that goes with this?

OpenStudy (imstuck):

Don't tell me, let me guess. You have to find the integers! right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Let the first number be x. So the next even number is x + 2. So their sum is x + (x+2) = 814 => 2x + 2 = 814 => 2x = 812 => x = 406 So the first number is 406 and the next numbr is 406+2 = 408

OpenStudy (anonymous):

So your numbers are 406 and 408. :)

OpenStudy (anonymous):

thank you so much now i understand how to do it

OpenStudy (anonymous):

You're welcome :)